please find the integral for both problems
s (x3 + 8x - 9 + 7 ) dx 2 [(x2 +1)9 5xdx
Answer #1

na 10 S (x + For 9+ 4) de + 20%2_ qu + logel + C : ( 2 ) +6 (120) Suda ; put 2+1=t andr-dt & non-alt Sedte ette6+)40 2 1 (+9

Please like the answer.
Answer #2

SOLUTION :



∫ (x^3 + x^(1/2) - 9 + 1/x) dx 

= 1/4  x^4  + 1/(3/2) x^(3/2) - 9x + ln(x) + C

=  x^4  / 4 + 2/3 x √x - 9x + ln(x) + C (ANSWER).


∫ (x^2 + 1)^9 * 5x dx

Let x^2 + 1 = u

=> 2x dx = du 

=> x dx = 1/2 du


So, given integral = ∫ u^9 * 5/2 du 

= u^10  / 10 * 5/2 + C

= u^10 / 4  + C

= (x^2 + 1)^10  / 4 + C (ANSWER).

answered by: Tulsiram Garg